AnimeGhoul8863:

LAST ONE! You are having a meeting with the CEO of a technology company. You have interpreted the number of laptops produced versus profit as the function P(x) = x4 − 3x3 − 8x2 + 12x + 16. Describe to the CEO what the graph looks like and, in general, how to sketch the graph without using technology. Use complete sentences, and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0).

AnimeGhoul8863:

@Mehek

Mehek:

You do the same thing where you factored or used the quadratic formula to solve for x and then found the y-intercept

AnimeGhoul8863:

If you factored this equation it would be P(x)=(x+2)(x-2)(x-4)(x+1) right?

Mehek:

Correct

NonportritV2:

Solve it step by step: \[x^4-3x^3-8x^2+12x+16=0\] Factor \[(x+1)(x-2)(x+2)(x-4)=0\] \[x+1=0 \] \[x=-1\] \[x - 2 =0\] \[x=2\] \[x-4 = 0\]

Mehek:

So when you solve for each x, you get your x-intercepts That's the first thing you'd plot

NonportritV2:

Unfortunately I don't know how to add graphs.

AnimeGhoul8863:

wait slow down a bit i just got confused so

AnimeGhoul8863:

once you factor this P(x) = x4 − 3x3 − 8x2 + 12x + 16.

Mehek:

So you got the factored part

AnimeGhoul8863:

you getP(x)=(x+2)(x-2)(x-4)(x+1)

Mehek:

Correct Now you need to know what each x is equal to

Mehek:

In order to do that, set each of those to 0 (x + 2) = 0 (x - 2) = 0 (x - 4) = 0 (x + 1) = 0

Mehek:

And then you solve for every one of them For the first one: x + 2 = 0 x + 2 - 2 = 0 - 2 x = -2 So one x-intercept is at (-2, 0) Do that for the rest of them

AnimeGhoul8863:

if im thinking correctly X intercepts (-2,0), (2,0), (4,0), (-1,0)

Mehek:

Yep

AnimeGhoul8863:

ok now what once we have the x intercepts what do we do now?

AnimeGhoul8863:

@Mehek

AnimeGhoul8863:

do i use Quadratic formula

AnimeGhoul8863:

if so then x=-2,2,4,-1

AnimeGhoul8863:

but i dont think thats what we need

AnimeGhoul8863:

@Mehek

AnimeGhoul8863:

@Shadow could you help me since she left T~T

Mehek:

You need the x-intercepts to graph

AnimeGhoul8863:

i already finished sorry i couldnt wait any longer